# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

### v2

@@ IV EXPERIMENT RESULTS
Botanic1 Garden Livox AVIA 1005_05 (566.8m) 0.773 1.053 0.640 7.931 0.917
1006_03 (686.2m) 1.796 2.075 3.017 6.527 2.471
1008_01 (855.9m) 1.152 1.577 14.829 4.445 1.590
VLP16 1005_05 (566.8m) 0.175 0.244 0.565 37.425 0.750
1006_03 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
1008_01 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
@@
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the time efficiency of system, and the validity of convergence criteria.
Figure 4: Our handheld sensor suite used in self-collected dataset, including the solid-state LiDAR Livox Avia with the built-in BMI088 IMU, and three other cameras.
IV-A Experiment Dataset
IV-B Odometry Accuracy Evaluation
@@
In odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2, DLIO, SLICT and Point-LIO. Some other frameworks are designed for spinning LiDAR hence they’re not included in comparison. All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. For comparison, we try to fine-tune other methods’ parameters to achieve their best performance. Besides, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
@@
In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Park_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians. These factors impose higher demands on the robustness of the LIO system. For the BotanicGarden datasets, we test both the solid-state (Livox AVIA) and the spinning (VLP16) LiDAR. Since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating long-duration drift under regular motion conditions. However, the abundance of unstructured scenes in the garden (e.g., thick woods, bushes) poses challenges for constructing residual constraints. In our self-collected datasets, we handheld the sensor suite and completed multiple closed-loop paths in different scenarios. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
@@
Concrete experiment results are demonstrated in Table II and Fig. 5. Our approach further optimizes the residual motion distortion via feedback mechanism based on the convergence criteria, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios. The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
TABLE III: Time Consumption Evaluation ($ms$)
@@
Min 4.65 3.15 3.43 21.33 11.95
Max 74.46 28.48 27.01 345.97 62.36
Mean 14.62 16.09 9.12 61.97 34.31
Median 11.71 16.68 8.50 51.49 34.98
Figure 6: The result of time efficiency evaluation. The cross line inside box denotes the median process time. Our AC-LIO achieves the second highest efficiency, only behind the lightweight DLIO framework. And the maximum iteration time also meets the real-time requirements.
IV-C Time Efficiency Evaluation
We calculate the runtime of each algorithm during iteration, with the results shown in Table III and Fig. 6. In terms of average and median processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the lightweight DLIO framework, and the maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input). Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
IV-D Convergence Criteria Validity Evaluation
TABLE IV: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
00 0 $0.5\sigma_{t}$ $[\sigma_{t},2\sigma_{t}]$ $3\sigma_{t}$
Park_3.2
0.624 0.610 0.582 0.616
Subway_3.3
2.841 2.419 1.670 1.788
Garden_1006
1.901 1.709 1.796 1.892
Garden_1008
1.197 1.199 1.152 1.178
To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels. While keeping other parameters the same, we only adjust the threshold of convergence criteria. Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$. The specific experimental results are shown in Table IV.
  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}({z})=\sigma_{t}$)
  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy. Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation. However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
V Conclusion and Future Work
In this paper, we proposed AC-LIO, a novel asymptotically and consistently converging LIO framework. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system. The approach of guiding asymptotic compensation based on convergence criteria is designed to promote consistent convergence of pointcloud registration, further improving the odometry accuracy while ensuring system robustness. Numerous experiment results indicate that our framework’s average RMSE is reduced by about 30.4% over the second best result, achieving higher accuracy in long-term, large-scale localization and mapping.
Future Work: We notice the performance of back propagation depends on the guidance of convergence criteria to some extend. The introduction of convergence criteria is essentially an attempt to overcome the disruption of the initial error on the back propagation during the iteration update. In future work, we will continue to study the convergence assessment of odometry, the combination of asymptotic backpropagation with other strategies, and the perception and adaptation of degenerate scenarios, to further explore the error-analyzable and scenario-interpretable SLAM framework.

### v3

@@ IV EXPERIMENT RESULTS
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the validity of convergence criteria, and the time efficiency of system.
IV-A Experiment Dataset
Figure 4: Our handheld sensor suite for self-collected dataset, including solid-state LiDAR Livox Avia with built-in BMI088 IMU, and three other cameras.
@@ IV-A Experiment Dataset
Botanic1 Garden Livox AVIA Garden_1005 (566.8m) 0.773 1.053 0.640 7.931 0.917
Garden_1006 (686.2m) 1.796 2.075 3.017 6.527 2.471
Garden_1008 (855.9m) 1.152 1.577 14.829 4.445 1.590
VLP16 Garden_1005 (566.8m) 0.175 0.244 0.565 37.425 0.750
Garden_1006 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
Garden_1008 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
@@
IV-B Experiment Settings:
  • We compared AC-LIO with current sota LIO frameworks that employ various methods to mitigate motion distortion, rather than LO frameworks without IMU integration. Some frameworks are designed for spinning LiDAR hence they are not included in comparison. Among the LIO frameworks, FAST-LIO2 compensates for distortion via IMU prior trajectory before iterative estimation, DLIO applies a coarse-to-fine pointcloud deskewing approach, SLICT utilizes continuous-time state estimation to optimize the motion distortion, and Point-LIO is an extremely high frequency LIO with a point-by-point strategy to avoid motion distortion.
  • All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. As our AC-LIO is implemented on the basis of FAST-LIO2, we control the same parameters especially with FAST-LIO2 (e.g., imu_noise, voxel_size, max_iteration, point_filter_num) for fair comparison. And we make every effort to fine-tune the parameters of other methods to achieve their best performance.
  • Since our AC-LIO computes the convergence criteria to guide the asymptotic distortion compensation during iterative update, we fix the iteration times for every epoch (i.e., single LiDAR frame). In contrast, FAST-LIO2 jumps out the iteration as the update gradient belows a threshold, which may lead to an accuracy difference with our AC-LIO. Therefore, we conducted ablation studies specifically (Sec. IV-D) to validate the performance of selective intra-frame smoothing.
IV-C Odometry Accuracy Evaluation
Figure 6: Accuracy evaluation in Botanic Garden dataset. The figure shows the result of sequence (a) Garden_1006 and (b) Garden_1008. For each sequence, the left part presents the trajectory comparison, with the top 10% states aligned to the ground truth, and the right part displays the cumulative error curve of APE (m) along with the corresponding real-time process efficiency. The results suggest that our AC-LIO achieves lower cumulative error while maintaining relatively good time efficiency (Sec. IV-E). The average runtime of our method is only slightly higher than the FAST-LIO2 framework.
@@ IV-C Odometry Accuracy Evaluation
The odometry accuracy evaluation aims to demostrate the overall accuracy performance of selected LIO frameworks and our AC-LIO. In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Parking_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians. These factors impose higher demands on the robustness of the LIO system. For the BotanicGarden datasets, we test both the solid-state (Livox AVIA) and the spinning (VLP16) LiDAR. Since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating long-duration drift under regular motion conditions. However, the abundance of unstructured scenes in the garden (e.g., thick woods, bushes) poses challenges for constructing residual constraints. In our self-collected datasets, we handheld the sensor suite and completed multiple closed-loop paths in different scenarios. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
@@
Concrete experiment results are demonstrated in Table II, Fig. 5 and Fig. 6. Our approach further optimizes the residual motion distortion via selective intra-frame smoothing strategy, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios. The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
IV-D Convergence Criteria Evaluation
TABLE III: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
[…]
Garden_1005 (566.8m) AVIA 1.053 1.063 1.063 1.004 0.943 0.743 0.907 1.120
VLP16 0.244 0.244 0.241 0.226 0.175 0.193 0.208 0.233
Figure 7: Validity evaluation of convergence criteria. These two figures present the ablation experiments on convergence criteria using Garden_1005 sequence: (a) for Livox AVIA and (b) for VLP16, showing the generalization between different LiDAR sensors. The gray dashed line refers the accuracy of FAST-LIO2. As threshold $\eta$ increases from zero, the accuracy improves because more sufficiently converged states are allowed to perform backpropagation. However, as $\eta$ continues to increase, states with larger initial errors are also allowed to backpropagate, which gradually degrades the overall accuracy.
To evaluate the rationality and validity of the proposed convergence criteria, we conduct two type of experiments. Rationality: Since the modeling of convergence criteria involves certain simplification and assumption, we examine the rationality through comparing the actual APR values with the theoretical value of $\sigma_{t}$, as the APR during static phase can be regarded as a reference of $\sigma_{t}$ under ideal conditions. We collect a static-movement-static sequence with Livox AVIA. As shown in Fig. 8, the actual APR values during the static phases are very close to $\sigma_{t}$, indicating good consistency between theoretical derivation of $\sigma_{t}$ and actual APR results. Furthermore, the APR increases and becomes more volatile during the movement phase. Too large an APR reflects a higher risk of insufficient convergence. We introduce the convergence criteria to prevent the asymptotic distortion compensation from starting with too high an initial error.
Validity: We examine the accuracy of AC-LIO at different $\eta$ threshold levels for both AVIA and VLP16, and compare with the benchmark of FAST-LIO2. While keeping other parameters the same, we only adjust the threshold of convergence criteria. The experimental results are shown in Table III and Fig. 7.
  • $\eta\in[0,0.5\sigma_{t}]$: No or rare backpropagation and asymptotic compensation is performed under this condition, reflecting the basic odometry accuracy.
  • $\eta\in[2.5\sigma_{t},3\sigma_{t}]$: Too high a threshold may lead to the backpropagation in case of large initial errors, potentially degrade the overall system accuracy and robustness.
Figure 8: Rationality evaluation of convergence criteria. The orange curve refers the average point-to-plane residual (APR) during a static-movement-static sequence in an indoor scenario. The APR values under static phases are very closed to the theoretical value of $\sigma_{t}$, indicating good consistency between theoretical derivation and actual APR results.
The results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ can achieve higher accuracy. As the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation. However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and gradually impacts odometry accuracy.
IV-E Time Efficiency Evaluation
We calculate the runtime of each algorithm during iteration, with the results shown in Table IV and Fig. 9. And the specific runtime curves can be found in Fig. 7. In terms of average processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the FAST-LIO2 framework and comparable to the lightweight DLIO framework. The maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input). Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
TABLE IV: Time Consumption Evaluation ($ms$)
@@ IV-E Time Efficiency Evaluation
Min 4.12 2.89 3.95 22.61 7.83
Max 70.69 66.38 37.22 460.24 50.84
Mean 8.55 7.18 8.57 48.62 24.48
Median 8.316 6.971 8.24 42.91 23.61
Figure 9: Time efficiency evaluation. The cross line inside box denotes the median process time. Our AC-LIO achieves the second highest efficiency, only behind the FAST-LIO2 and comparable to the lightweight DLIO. The maximum runtime also meets the real-time requirements.
V Conclusion
In this paper, we proposed AC-LIO, a novel discrete state LIO framework with selective intra-frame smoothing. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system. The approach of guiding asymptotic compensation based on convergence criteria is designed to further suppress the motion distortion within LiDAR frame, improving the odometry accuracy while ensuring system robustness. Extensive experiments indicate that our framework achieves an average RMSE reduction of approximately 30.4% over the second best method, demonstrating higher accuracy in long-term, large-scale localization and mapping. In terms of time efficiency, our framework ranks second place among the evaluated frameworks, just following the FAST-LIO2 and comparable to the lightweight DLIO, achieving a balance between accuracy and real-time performance.